# Automated Model Selection for Tabular Data

### v1

@@ 1 Introduction
Feature interactions occur when the combination of two or more features provides insights and information that cannot be derived from considering each feature in isolation. For example, the Adult Income dataset [Dua et al., 2019] contains features like education-level, marital-status, race, gender etc., which are not equally important to income and are related in ways that might in some combinations, strongly predict income when together than treated separately. For example, it is possible that when gender=female and marital-status=married and age $\in$ 25-35 the combination of these factors would indicate a higher likelihood of child-bearing and maternity leave which would have a significant effect on one’s income compared to combining these indicators’ contribution individually: because of the way the features interact. Thus, in this context of classification and regression on such tabular data, the choice of features used plays a pivotal role in the performance of our predictors. The role of feature interactions within tabular datasets has received increasing attention in recent years in works like [Xie et al., 2020] and [Liu et al., 2020]. Linear Regression (LR) models linearly aggregate features without modeling complex feature interactions. This can limit performance when interdependencies exist. Recognizing and effectively modeling feature interactions is pivotal in enhancing the overall performance and interpretability of models.
@@ 3.2 Objective and Framework
Figure 1: Feature importances in the real-life Adult Income dataset.
We see the potential for automating this search for an optimum model, while keeping computational costs small. We will optimise our search using prior assumptions on the extent of feature interactions and using feature importances to cut our search space short. At the milestone stage, we have implemented priors on feature interactions, and in next steps will incorporate prioritising individual feature importances in our search.
We plan to optimize the metrics of the model such as $R^{2}$ value (coefficient of determination), MSE (mean squared error), Log-likelihood and AIC (Akaike Information Criterion) to show the performance of our models. We also track runtime of feature search to evaluate computational efficiency of the search. The main libraries we use are from various sklearn, numpy, statsmodels and pandas functionalities. We work with linear regression models applied on categorical variables. For this we apply sklearn’s LinearRegression functionality upon various one-hot encoded Python pandas dataframes. We also plan to explore generalized models and ensemble models. We encode categorical features using one-hot encoding before fitting linear regression models.
@@ 4.2 Synthetic data generation
Based on the above analysis on the Adult dataset, we fell back to creating a synthetic dataset where we incorporate some desired feature interactions. These feature interactions serve as the ground truth to evaluate our method on. We generated a synthetic dataset we call ‘Adult2’, deliberately incorporating feature interactions. We assigned weights to the features and combinations of features as illustrated in figure 4. These weights were then used to combine the contributions of the features into the target (label) attribute, salary.
@@
We generate data attributes by uniform random generation over the possible value space. In our setting, we introduced interactions via a designed feature, ‘MaritalStatusGender’ to capture the interaction between ‘MaritalStatus’ and ‘Gender’. We set the weights for this feature so that it is not linear in terms of the individual components (see fig. 4). This feature, along with a subset of base features is weighted and added up to generate our target attribute ‘Salary’. This process is illustrated in figure 2. Let all the features used be called ($F_{used}$). At the end, to introduce variability, we inject a noise factor to Salary, creating a 2000 sample dataset. A snippet from this dataset is displayed in figure 3.
@@ 5.1 Priority-based random grid search
Initially, we considered a full grid search, exploring nearly $2^{n+{n\choose 2}+{n\choose 3}+...+{n\choose n}}$ potential combinations of features and interactions, where n represents the number of features in the dataset, and (${n\choose r}$ represents the interactions of degree r. Due to the computational complexity and substantial computation time associated with all these possibilities, we decided to restrict to exploring interactions only via the second order. This still leaves $2^{n+{n\choose 2}}$ possible feature subsets involving the base features and the 1st order interaction features. So we had to simplify it further by assuming that there would be very few feature interactions among the (${n\choose 2}$ combinations involved in the true model. We took the number of 2nd order feature interaction terms as being [0, 1, 2].
@@
9:     $F_{\text{cur}}=F_{\text{sampled}}\cup I_{\text{selected}}$
@@
4:Fit a linear model with ($F_{\text{selected}},y$) ans store p-values and std error for all features
@@ 5.2 Greedy search
Figure 5: Evolution of the best feature set over iterations of the priority search algorithm. From the figure, we observe that the true model is found after 3483 iterations
@@ 6.1 Priority-based random grid search
Figure 5 shows our result for priority search. We observe that the true model is captured at iteration 3434. The better scoring model found later is overfit with more parameters (a superset of the already found true model), both have the almost same score. We also achieve full 99.6% accuracy (because the synthetic data was made using the same linear additions we capture in our LR models).
The following graph (Figure 6) depicts the $R^{2}$ value of the best model identified in relation to the number of times the algorithm was executed. On average, the peak performance of the best model is observed around the $3434^{th}$ iteration. Total iterations for all possible feature subsets for our dataset (under highest degree of interaction being 2) would’ve been $2^{8+{8\choose 2}}$.
Figure 6: Best performance achieved ($R^{2}$) vs N with the priority search algorithm.
@@ 6.2 Greedy Search
Considering the 8 base features in the synthetic data (Figure 3), and the 2nd order interactions features (28), there are 36 features and the potential feature space after one-hot encoding is 953 (coefficients). This results in a total of $2^{953}$ feature subsets, making it computationally intensive to find the best model. Thus we use greedy approach to optimize the search space. Figure 7, 8 shows the learning curve for Forward selection algorithm and backward elimination respectively considering the metrics $R^{2}$, MSE, AIC, Log-likelihood.
@@ 6.3 Discussion
Table 1 summarizes the results from priority search, backward elimination and forward selection. From the table 1, we can conclude that the optimal solution using forward selection and backward elimination may differ.
@@
    Random Priority search achieved a 99.57% $R^{2}$ with an average runtime of 2 to 3 minutes, capturing a near-optimal (near true) model. This method required fewer iterations and took less time.
@@
    The Forward selection achieved a 99.62% $R^{2}$ on the synthetic data with an average runtime of 20 to 30 minutes, capturing the true model.
@@
    The Backward Selection achieved over 99.68% $R^{2}$ with an average runtime of 10 to 15 minutes, capturing a near-optimal solution.
From the experiments, the forward selection seems to give us the true model when compared to the other approaches. Priority based random grid search required the fewer iterations and runtime when compared with the other greedy approaches. This highlights a tradeoff between computational efficiency and performance. Importantly, all methods outperformed baseline linear models, demonstrating the value of modeling feature interactions.
@@
Table 1: Comparing priority search, backward elimination and forward selection statistics. We also report the generated model against the true model, which is S $\sim$ C + E + M*G + O + W.
@@ 7 Conclusion
In this work, we investigated automated model selection incorporating feature interactions for tabular datasets. We hypothesized that combinations of features can be more predictive than individual contributions. We proposed three approaches: a Priority-based Random Grid Search, a Greedy Search method with forward selection, and backward elimination. The Priority search uses priors to efficiently explore feature subsets. The Greedy approach iteratively adds/removes features based on impact. Experiments conducted on real world Adult dataset reveals that there are no singificant interactions hence we fall back to create synthetic dataset on which forward selection algorithms achieves true model whereas priority-based random grid search achieves near optimal model with minimal computational time.
Figure 8: Backward Elimination performance. We terminate backward elimination due to sudden surge in AIC, which implies that a crucial features has been removed.
@@ 8 Future work
The first area of improvement involves refining the Priority based Random Grid search by incorporating a learned prior probability distribution on k. Additionally, expanding the size of k and implementing downweighting based on feature size is planned, with the aim of addressing large feature spaces such as country x occupation, which may result in a substantial number of possible values. Another avenue for improvement is the optimization of the Greedy algorithm by considering H-statistics. The exploration of various regularization methods for feature selection is also slated for investigation. Lastly, a comparative SHAP analysis between forward selection and backward elimination will be conducted to gain insights into the interpretability and impact of these feature selection approaches. These proposed enhancements collectively aim to advance the effectiveness and efficiency of the current research methodology.

### v2

@@ 1 Introduction
Feature interactions occur when the combination of two or more features provides insights and information that cannot be derived from considering each feature in isolation. For example, the Adult Income dataset [Dua et al., 2019] contains features like education-level, marital-status, race, gender etc., which are not equally important to income and are related in ways that might in some combinations, strongly predict income when together than treated separately. For example, it is possible that when gender=female and marital-status=married and age $\in$ 25-35 the combination of these factors would indicate a higher likelihood of child-bearing and maternity leave which would have a significant effect on one’s income compared to combining these indicators’ contribution individually: because of the way the features interact. Thus, in this context of classification and regression on such tabular data, the choice of features used plays a pivotal role in the performance of our predictors. The role of feature interactions within tabular datasets has received increasing attention in recent years in works like [Xie et al., 2020] and [Liu et al., 2020]. Linear Regression (LR) models linearly aggregate features without modeling complex feature interactions. This can limit performance when inter-dependencies exist. Recognizing and effectively modeling feature interactions is pivotal in enhancing the overall performance and interpretability of models.
@@ 3.2 Objective and Framework
Figure 1: Feature importance’s in the real-life Adult Income dataset.
We see the potential for automating this search for an optimum model, while keeping computational costs small. We optimise our search using prior assumptions on the extent of feature interactions and using feature importance’s to cut our search space short. We optimize the metrics of the model such as $R^{2}$ value (coefficient of determination), MSE (mean squared error), Log-likelihood and AIC (Akaike Information Criterion) to show the performance of our models. In addition, we track runtime of feature search to evaluate computational efficiency of the search. We use the following frameworks for out experiments: sklearn, numpy, statsmodels OLS and pandas. Since, we work with linear regression models applied on categorical variables, we apply sklearn’s LinearRegression functionality upon various one-hot encoded pandas dataframes. We encode categorical features using one-hot encoding before fitting linear regression models.
@@ 4.2 Synthetic data generation
Based on the above analysis on the Adult dataset, we aim to create a synthetic dataset where we incorporate some desired feature interactions. These feature interactions serve as the ground truth to evaluate our method on. We generated a synthetic dataset we call ‘Adult2’, deliberately incorporating feature interactions. We assigned weights to the features and combinations of features as illustrated in figure 4. These weights were then used to combine the contributions of the features into the target (label) attribute, salary.
@@
We generate data attributes by uniform random generation over the possible value space. In our setting, we introduced interactions via a designed feature, ‘MaritalStatusGender’ to capture the interaction between ‘MaritalStatus’ and ‘Gender’. We set the weights for this feature so that it is not linear in terms of the individual components (see fig. 4). This feature, along with a subset of base features is weighted and added up to generate our target attribute ‘Salary’. This process is illustrated in figure 2. Let all the features used be called ($F_{used}$). At the end, to introduce variability, we inject a noise factor to Salary, creating a dataset with 2000 samples. A snippet from this dataset is displayed in figure 3.
@@ 5 Techniques and methods
5.1 Computational complexity of the feature space
Initially, we considered a full grid search, exploring nearly $2^{n+{n\choose 2}+{n\choose 3}+...+{n\choose n}}$ potential combinations of features and interactions, where n represents the number of features in the dataset, and (${n\choose r}$ represents the interactions of degree r. Due to the computational complexity and substantial computation time associated with all these possibilities, we decided to restrict to exploring interactions only via the second order. This still leaves $2^{n+{n\choose 2}}$ possible feature subsets involving the base features and the 1st order interaction features.
@@ 5.2 Priority-based random grid search
So we had to simplify it further by assuming that there would be very few feature interactions among the (${n\choose 2}$ combinations involved in the true model. We took the number of 2nd order feature interaction terms as being [0, 1, 2].
@@
9:     $F_{\text{cur}}=F_{\text{sampled}}\cup I_{\text{selected}}$ (to remove redundant features)
@@
4:Fit a linear model with ($F_{\text{selected}},y$), store p-values and std error for all features
@@ 5.3 Greedy search
Figure 5: Evolution of the best feature set over iterations of the priority search algorithm. From the figure, we observe that the algorithm reached near optimal model
@@ 6.1 Priority-based random grid search
Figure 5 shows our result for priority search. We observe that the algorithm achieved the near optimal model ($\sim A+C*R+E+G*M+O+W$) after 2766 iterations. We also achieve an optimal $R^{2}$ of 0.996 because the synthetic data was made using the same linear additions we capture in our LR models.
Figure 6 depicts the $R^{2}$, MSE, log likelyhood and AIC value of the best model identified in relation to the number of times the algorithm was executed. On average, the peak performance of the best model is observed around the $100^{th}$ iteration. Total iterations for all possible feature subsets for our dataset (under highest degree of interaction being 2) would’ve been $2^{8+{8\choose 2}}$.
Figure 6: Priority search performance. The x-axis employs a logarithmic scale.
@@ 6.2 Greedy Search
Considering the 8 base features in the synthetic data (Figure 3), and the 2nd order interactions features (28), there are 36 features and the potential feature space after one-hot encoding is 953 (coefficients). This results in a total of $2^{953}$ feature subsets, making it computationally intensive to find the best model. Thus we use greedy approach to optimize the search space. Figure 7, 8 shows the learning curve for Forward selection algorithm and backward elimination respectively considering the metric.
@@ 6.3 Discussion
Table 1 summarizes the results from priority search, backward elimination and forward selection. From the table 1, we can conclude that the optimal solution using priority-based grid search, forward selection and backward elimination may differ.
@@
    Random Priority search achieved a 0.9966 $R^{2}$ with an average runtime of 2 to 3 minutes, capturing a near-optimal (near true) model. This method required fewer iterations and took less time.
@@
    The Forward selection achieved a 0.9962 $R^{2}$ on the synthetic data with an average runtime of 20 to 30 minutes, capturing the true model.
@@
    The Backward Selection achieved over 0.9968 $R^{2}$ with an average runtime of 10 to 15 minutes, capturing a near-optimal solution.
From the experiments, the forward selection seems to give us the true model when compared to the other approaches. Priority based random grid search required the fewer computational runtime when compared with the other greedy approaches. This highlights a trade-off between computational efficiency and performance. Importantly, all methods outperformed baseline linear models (models without interactions), demonstrating the value of modeling feature interactions.
@@
Table 1: Comparing priority search, backward elimination and forward selection statistics. We also report the generated model against the true model, which is S $\sim$ C + E + M*G + O + W. We use google colab with T4 GPU runtime with 12GB RAM to run the experiments.
@@ 7 Conclusion
In this work, we investigated automated model selection incorporating feature interactions for tabular datasets. We hypothesized that combinations of features can be more predictive than individual contributions. We proposed three approaches: a Priority-based Random Grid Search, a Greedy Search method with forward selection, and backward elimination. The Priority search uses priors to efficiently explore feature subsets. The Greedy approach iteratively adds/removes features based on impact. Experiments conducted on real world Adult dataset reveals that there are no significant interactions hence we fall back to create synthetic dataset on which forward selection algorithms achieves true model whereas priority-based random grid search achieves near optimal model with minimal computational time.
Figure 8: Backward Elimination performance. We terminate backward elimination due to sudden surge in AIC, which implies that a crucial features has been removed. Debugging this revealed that the removed feature is $M*G$ which is a crucial interaction feature in the true model. Hence, we backtrack one step to restore this feature and terminate the algorithm.
@@ 8 Future work
In future work, there are several areas for enhancement in our research methodology. Firstly, we plan to refine the Priority-based Random Grid search by integrating a learned prior probability distribution on parameter k. Additionally, we intend to extend the range of parameter k and introduce down-weighting based on feature size to better accommodate large feature spaces such as country $\times$ occupation, which may contain numerous possible values. Finally, we aim to conduct a comparative SHAP analysis between forward selection and backward elimination to assess the interpretability and impact of these feature selection approaches. These proposed enhancements collectively seek to enhance the effectiveness and efficiency of our current research methodology.